### v2

# Ensembling Portfolio Strategies for Long-Term Investments: A Distribution-Free Preference Framework for Decision-Making and Algorithms

Abstract
This paper investigates the problem of ensembling multiple strategies for sequential portfolios to outperform individual strategies in terms of long-term wealth. Due to the uncertainty of strategies’ performances in the future market, which are often based on specific models and statistical assumptions, investors often mitigate risk and enhance robustness by combining multiple strategies, akin to common approaches in collective learning prediction. However, the absence of a distribution-free and consistent preference framework complicates decisions of combination due to the ambiguous objective. To address this gap, we introduce a novel framework for decision-making in combining strategies, irrespective of market conditions, by establishing the investor’s preference between decisions and then forming a clear objective. Through this framework, we propose a combinatorial strategy construction, free from statistical assumptions, for any scale of component strategies, even infinite, such that it meets the determined criterion. Finally, we test the proposed strategy along with its accelerated variant and some other multi-strategies. The numerical experiments show results in favor of the proposed strategies, albeit with small tradeoffs in their Sharpe ratios, in which their cumulative wealths eventually exceed those of the best component strategies while the accelerated strategy significantly improves performance.
Keywords: Online Learning, Multi-Strategy, Ensembling Strategies, Preference, Decision Making.
1 Introduction
In this paper, we investigate the problem of ensembling multiple sequential portfolio strategies such that the combinatorial strategy could eventually outperform all the component ones during a long-term investment. The component strategies may include a broad range of prediction and analysis-based strategies or even naive ones, such as the equal allocation strategy, that an investor can observe over time. A natural problem arises in that, due to the uncertainty of the future, the performances of the observed strategies are not guaranteed and seem uncertain. Moreover, many sophisticated strategies depend on a random market with specific statistical assumptions, which are often invalid in reality, as the future market could evolve much differently from the hypothetical models. Hence, an investor could tackle the problem by ensembling several strategies rather than believing in a single one, which is probably a well-known common approach for noise reduction in machine learning prediction, like the fusion models. Unfortunately, since there has not been a concept to measure the preference of choice between component strategies that do not rely on known distributions, which is required for the traditional expected utility framework, the distribution-free preference framework between the ensembles of collective strategies also does not exist. In order to deal with this challenge, we primarily put attention on well-forming a particular objective of ensembling strategies through establishing a preference framework for making decisions without relying on statistical assumptions on market data.
Given the significant interest and importance of collective learning methods and ensembles of trading or investment strategies in the literature, particularly within practical investment contexts, there has been a substantial volume of research, mostly experimental. To underscore our later arguments, we briefly reference several recent articles that follow the increasing popularity of algorithmic trading under the umbrella of fintech and machine learning such as Nti et al. (2020), Padhi et al. (2022), Yu et al. (2023), Sun et al. (2023) and Deng et al. (2024). The similarity in contemporary research is they only focus on prediction and lack a uniquely theoretical framework for the formulation of an economic agent’s objective and preference. In general, the problem of strategies ensemble can be equivalently cast as the problem of making a portfolio of assets’ allocation, which results in the inherent uncertainty in the decision of choice. Consequently, their performances in out-of-sample long-term investments are not secured. Moreover, as a particular framework for making an ensemble has not been established, the problem of an investor learning in a market with an infinitely large amount of agents’ strategies is difficult to be investigated. These issues motivate us to propose a framework for making decisions on ensembling various component strategies, with a clearly defined and consistent preference over the lifetime of a market.
The main contributions and organization of the paper. By convention, the primary objective of an investor in this paper is to create an ensemble of component strategies that eventually exceed all of them in terms of wealth at some future periods. In Section 2, we firstly establish a model of an investor in a market who sequentially decides combinations of various observable component strategies of no-short portfolios over time. Then, we introduce a general framework of distribution-free and time-consistent preference based on the cumulative wealth of a strategy, which constitutes the investor’s goal. Based on this theoretical framework, the criterion for creating an ensemble of strategies is specified. In Section 3, since the preferences between strategies are defined on each sequence of unknown market data, we propose a particular online learning strategy to determine combinations of component strategies, ensuring the investor’s preference is universally guaranteed for any possibilities of market data. In the case of a market with an infinitely large number of observable strategies, we propose a modified algorithm to reduce dimensionality by learning on a smaller set of representative strategies, thus ensuring universal preference.
In Section 4, we conduct experiments to numerically approximate the proposed strategy and explore further accelerated variants by leveraging the flexibility of the original strategy’s mechanism. The proposed strategies are tested and compared with other online learning strategies, serving as alternative methods for ensemble multi-strategies, in terms of empirical performances such as the Sharpe ratio, average growth rate, and final cumulative wealth. The dataset utilized for these experiments is sourced from The Center for Research in Security Prices, LLC, spanning a comprehensive range of 27 years, which includes many historically significant market events and comprises six blue-chip stock symbols. Various experimental scenarios are conducted for both small and large scales of component strategies, involving sequential Mean-Variance portfolios with different risk profiles. In all experiments, while the other multi-strategies often performed poorly, the wealth generated by the proposed strategy eventually exceeds that of the best component ones, albeit with slight tradeoffs in the Sharpe ratios. Moreover, the accelerated strategy also significantly increases cumulative wealth compared with the original one, though it is not theoretically guaranteed but rather heuristic.
2 Model settings and formalizations
[…]
At each $n$, let $b_{n}:\mathbb{R}_{++}^{m\times(n-1)}\to\mathcal{B}^{m}$ denote the choice of portfolio based on $x_{1}^{n-1}$, and let the strategy associated with these portfolios be denoted by the infinite sequence $\big{(}b_{n}\big{)}\coloneqq\left\{b_{n}\right\}_{n=1}^{\infty}\in\mathcal{B}^{m\times\infty}$; accordingly, a constant strategy of a fixed portfolio $b$ is denoted as solely $\big{(}b\big{)}$, without a timing index. Given the initial capital $S_{0}\big{(}b_{0}\big{)}\eqqcolon S_{0}=1$, assuming there are no commission fees, the cumulative wealth and its corresponding exponential average growth rate after $n$ periods of investment using a generic strategy $\left(b_{n}\right)$ are defined respectively as:
$S_{n}\left(b_{n}\right)\coloneqq{\displaystyle{\displaystyle\prod_{i=1}^{n}}\left\langle b_{i},x_{i}\right\rangle}\text{ and }W_{n}\left(b_{n}\right)\coloneqq\dfrac{1}{n}\log S_{n}\left(b_{n}\right).$
By employing the definitions of cumulative wealth and growth rate, we propose a conceptual framework that generalizes the evolution of a portfolio strategy in long-term investment, as defined in Definition 1 below, and also provide an illustration in Example 1.
Definition 1.
A strategy $\big{(}\hat{b}_{n}\big{)}$ is said to almost always exceed another strategy $\big{(}\bar{b}_{n}\big{)}$ if there exists $M>1$ such that $S_{n}\big{(}\hat{b}_{n}\big{)}>S_{n}\big{(}\bar{b}_{n}\big{)}$ for all $n\geq M$ (the term “almost” can be omitted for $M=1$).
As the negation of the Definition 1, the strategy $\big{(}\hat{b}_{n}\big{)}$ does not always exceed the strategy $\big{(}\bar{b}_{n}\big{)}$ if for any $M\geq 1$, there exists $n_{M}\geq M$ such that $S_{n_{M}}\big{(}\hat{b}_{n_{M}}\big{)}\leq S_{n_{M}}\big{(}\bar{b}_{n_{M}}\big{)}$; moreover, if there are infinitely many $n_{M}$ such that the strict inequality holds, the strategy $\big{(}\hat{b}_{n}\big{)}$ can be said to be infinitely often exceeded by the strategy $\big{(}\bar{b}_{n}\big{)}$. This way of defining the relative performance in terms of the growth of strategies is general in the sense that it does not rely on any specific statistical properties of the assets’ returns. Hence, the cumulative wealth of a strategy will either be almost always exceeded or not by that of another, regardless of the behavior of the infinite sequence $x_{1}^{\infty}$.
Example 1.
Figure 1 illustrates an example of the evolutions of some strategies taken from the experiment section. The left graphic shows two strategies often exceeding each other, while the right graphic shows a strategy almost always exceeding the remainder, except for a few observations around time $n=3400$, where they are temporarily equivalent.
Figure 1: Evolutions of component M-V strategies over time with different levels of risk aversion.
The main point of introducing the evolutionary framework is that a strategy which is not always exceeded by other strategies will eventually exceed them after a finite number of time periods from any observation point. However, although using Definition 1 alone is able to distinguish any pairs of strategies, it seems not convenient for formalizing a preference ranking essential for constructing an ensemble. In the next sections, we propose an alternative but closely related preference relation framework which constitutes a criterion for strategy combination corresponding to the investor’s objective. Then, we demonstrate its link to the evolution of cumulative wealth of strategy, wherein a strategy cannot always exceed another one that is strictly preferred.
2.1 Preference relation and combination criterion
Motivated by the fact that the amount of market data is infinitely growing rather than fixed and wholly available, we need a preference relation that demonstrates a stable ranking over time to compare any two strategies, which implies that the relation does not change as the assets’ returns are gradually observed as time evolves. For this reason, such a preference should be defined over all the strategies in a single infinite-dimensional space $\mathcal{B}^{m\times\infty}$. The following Definition 2 introduces an appropriate preference in an asymptotic sense, posing consistency for the ranking.
Definition 2.
Given an infinite sequence of assets’ returns $x_{1}^{\infty}$ and two generic strategies $\big{(}\hat{b}_{n}\big{)}$, $\big{(}\bar{b}_{n}\big{)}$, the preference relations along with their corresponding notations are defined as follows:
  1. 1.
    $\big{(}\bar{b}_{n}\big{)}\sim\big{(}\hat{b}_{n}\big{)}$: Neither one of them is preferred to the other, or they are indifferent, as
    $\big{(}\bar{b}_{n}\big{)}\sim\big{(}\hat{b}_{n}\big{)}\Leftrightarrow\lim_{n\to\infty}\big{(}W_{n}\big{(}\hat{b}_{n}\big{)}-W_{n}\big{(}\bar{b}_{n}\big{)}\big{)}=0.$
  2. 2.
    $\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}$: $\big{(}\bar{b}_{n}\big{)}$ is weakly preferred to, or at least as good as, $\big{(}\hat{b}_{n}\big{)}$, as
    $\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}\Leftrightarrow\limsup_{n\to\infty}\big{(}W_{n}\big{(}\hat{b}_{n}\big{)}-W_{n}\big{(}\bar{b}_{n}\big{)}\big{)}\leq 0.$
  3. 3.
    $\big{(}\bar{b}_{n}\big{)}\succ\big{(}\hat{b}_{n}\big{)}$: $\big{(}\bar{b}_{n}\big{)}$ is strictly preferred to, or better than, $\big{(}\hat{b}_{n}\big{)}$, as
    $\big{(}\bar{b}_{n}\big{)}\succ\big{(}\hat{b}_{n}\big{)}\Leftrightarrow\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}\land\big{(}\bar{b}_{n}\big{)}\nsim\big{(}\hat{b}_{n}\big{)}.$
Hence, the subsequent logic deduction follows the definition of relations as follows:
$\big{(}\bar{b}_{n}\big{)}\succ\big{(}\hat{b}_{n}\big{)}\veebar\big{(}\bar{b}_{n}\big{)}\sim\big{(}\hat{b}_{n}\big{)}\Leftrightarrow\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}\text{ and }\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}\land\big{(}\hat{b}_{n}\big{)}\succsim\big{(}\bar{b}_{n}\big{)}\Leftrightarrow\big{(}\bar{b}_{n}\big{)}\sim\big{(}\hat{b}_{n}\big{)}.$
The preference relations defined above provide a consistent ranking for the investor, who prefers cumulative wealth, to evaluate utility for any strategy through its equivalent performance metric, the average growth rate. In an asymptotic sense with a timing factor, two similarly preferred strategies might not have the same growth rate all the time, but their difference is negligible as time evolves. This is more general than the usual sense of indifference in preference without a timing factor. Akin to a conventional preference relation, the proposed preference relation also possesses the following properties of a rational ranking, according to the fundamental axioms.
Proposition 1.
For any infinite sequence of assets’ returns $x_{1}^{\infty}$, the preference relations established in Definition 2 satisfy the following properties:
  1. 1.
    Completeness: $\big{(}\hat{b}_{n}\big{)}\succsim\big{(}\bar{b}_{n}\big{)}\lor\big{(}\bar{b}_{n}\succsim\big{(}\hat{b}_{n}\big{)}$, $\forall\big{(}\hat{b}_{n}\big{)},\big{(}\bar{b}_{n}\big{)}\in\mathcal{B}^{m\times\infty}.$
  2. 2.
    Reflexivity: $\big{(}\hat{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}$, $\forall\big{(}\hat{b}_{n}\big{)}\in\mathcal{B}^{m\times\infty}$.
  3. 3.
    Transitivity: $\big{(}\hat{b}_{n}\big{)}\succsim\big{(}\bar{b}_{n}\big{)}\land\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\tilde{b}_{n}\big{)}\Rightarrow\big{(}\hat{b}_{n}\big{)}\succsim\big{(}\tilde{b}_{n}\big{)}$, $\forall\big{(}\hat{b}_{n}\big{)},\big{(}\bar{b}_{n}\big{)},\big{(}\tilde{b}_{n}\big{)}\in\mathcal{B}^{m\times\infty}$.
  4. 4.
    Asymmetry: $\big{(}\hat{b}_{n}\big{)}\succ\big{(}\bar{b}_{n}\big{)}\Rightarrow\big{(}\bar{b}_{n}\big{)}\not\succsim\big{(}\hat{b}_{n}\big{)}$, $\forall\big{(}\hat{b}_{n}\big{)},\big{(}\bar{b}_{n}\big{)}\in\mathcal{B}^{m\times\infty}$.
Proof.
All properties could be checked straightforwardly using Definition 2. ∎
Given a set of strategies, since the infinite sequence $x_{1}^{\infty}$ is unknown, their evolutions of cumulative wealths are not definitively determined at any time. Thus, any competitive strategy can only guarantee to be at least as good as the best strategy at some future particular observation periods. In order to create such a competitive strategy conveniently and comprehensively, it is necessary to define a single baseline strategy against which it will compete. This baseline strategy, representing the best possible cumulative wealth among the observed strategies, is formulated based on the realizations of assets’ returns during the respective time periods.
Definition 3.
(Baseline strategy). For an infinite sequence of assets’ returns $x_{1}^{\infty}$ and a set of strategies $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\coloneqq\big{\{}1,\ldots,k\big{\}}\big{\}}$, the baseline strategy, denoted as $\big{(}\ddot{b}_{n}\big{)}$ hereafter, is defined as the best merged strategy of all the component ones, where $S_{n}\big{(}\ddot{b}_{n}\big{)}\coloneqq\max_{\alpha\in[k]}S_{n}\big{(}b_{n}^{\alpha}\big{)}$ for all $n$.
By definition, the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$ is the envelope of all cumulative wealth of observed strategies over time, so $\big{(}\ddot{b}_{n}\big{)}\succsim\big{(}b_{n}^{\alpha}\big{)}$ for all $\alpha\in\left[k\right]$. Since the construction of the baseline strategy relies on future data of assets’ returns at every time period, it is inherently inaccessible to the investor. However, the investor could instead create a multi-strategy (or an ensemble), denoted as $\big{(}b_{n}\big{)}$ and termed the combinatorial strategy henceforth, to compete with the baseline strategy. Despite the inaccessibility of the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$, it facilitates us to evaluate the preference between the combinatorial strategy and a single consistent strategy in the infinite-dimensional space $\mathcal{B}^{m\times\infty}$, thereby ensuring that the investor’s objective is met. To clarify this point, it is simple but worth noting that if $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$, then $\big{(}b_{n}\big{)}\succ\big{(}b_{n}^{\alpha}\big{)}$ for all $\alpha\in\left[k\right]$, due to the transitivity and asymmetry of preference, but the reverse is not necessarily true. Since strict preference is favored over a weak one, let us establish a specific criterion for combining the strategies.
Proposition 2.
(Combination criterion). Given an infinite sequence of assets’ returns $x_{1}^{\infty}$, the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$ cannot always exceed a combinatorial strategy $\big{(}b_{n}\big{)}$ if $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$.
Proof.
See Appendix. ∎
Remark.
The reverse reasoning of the Combination criterion is not correct since the baseline strategy could always exceed the combinatorial strategy over time despite $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$. For instance:
$1<S_{n}\big{(}\ddot{b}_{n}\big{)}\big{/}S_{n}\big{(}b_{n}\big{)}<2,\forall n\Rightarrow\big{(}b_{n}\big{)}\sim\big{(}\ddot{b}_{n}\big{)}\Rightarrow\big{(}b_{n}\big{)}\nsucc\big{(}\ddot{b}_{n}\big{)},$
due to the asymmetry of preference, so the weak preference is not sufficient for a combination.
According to the Combination criterion, the strict preference $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ implies that the cumulative wealth of the combinatorial strategy will eventually match or even surpass all those of the component strategies simultaneously after a finite number of time periods. Notably, it also includes the possibility of the combinatorial strategy almost always exceeding the baseline strategy. Furthermore, this guarantee does not rely on any statistical assumptions and is also not biased toward the selection of the starting point of observations [^1] In the literature, investment strategies and algorithms for ensembling them are commonly tested on a specific dataset, and the strategy with the highest final wealth is considered the best one. However, the choice of a dataset for out-of-sample testing can be biased towards certain strategies and does not account for longer horizons. Therefore, the Combination criterion generalizes a target for ensembling the component strategies. By following a combinatorial strategy that is not always exceeded by all other strategies concurrently, the investor can be confident of eventually accumulating the highest final wealth at certain points in the future.. Additionally, as an example to reiterate the significance and convenience of the baseline strategy as mentioned above, it is easy to see that creating a combinatorial strategy such that $\big{(}b_{n}\big{)}\succ\big{(}b_{n}^{\alpha}\big{)}$ for all $\alpha\in\left[k\right]$ does not guarantee that it will exceed all component strategies simultaneously.
The universality of a preference under unknown future assets’ returns. Since the preferences of strategies depend on an unknown specific sequence $x_{1}^{\infty}$, the following Claim 1 formalizes the universality of a preference over all possibilities of $x_{1}^{\infty}$ in order to decide a choice between two strategies. Generally, a weak preference is more universal than the reverse one if the strict relation holds for some possibilities of $x_{1}^{\infty}$ while the indifferent relation holds for the remaining others.
Claim 1.
Consider two generic strategies $\big{(}\hat{b}_{n}\big{)}$ and $\big{(}\bar{b}_{n}\big{)}$ and the infinite-dimensional distribution space $\mathbb{R}_{++}^{m\times\infty}$ of all instances of the sequence $x_{1}^{\infty}$. By the completeness and asymmetry of preference, if $\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}$ everywhere in $\mathbb{R}_{++}^{m\times\infty}$ while $\big{(}\hat{b}_{n}\big{)}\succsim\big{(}\bar{b}_{n}\big{)}$ only in a non-empty set $D\subset\mathbb{R}_{++}^{m\times\infty}$, then $\big{(}\bar{b}_{n}\big{)}\succ\big{(}\hat{b}_{n}\big{)}$ in $\mathbb{R}_{++}^{m\times\infty}\backslash D$. In this context, an economically rational investor in the market universally prefers the strategy $\big{(}\bar{b}_{n}\big{)}$ over the strategy $\big{(}\hat{b}_{n}\big{)}$, or the preference $\big{(}\bar{b}_{n}\big{)}\succsim\big{(}\hat{b}_{n}\big{)}$ is said to be more universal than the preference $\big{(}\hat{b}_{n}\big{)}\succsim\big{(}\bar{b}_{n}\big{)}$. This is reasonable by the fact that the possibility of the strategy $\big{(}\hat{b}_{n}\big{)}$ being strictly preferred does not exist. In the case $\big{(}\bar{b}_{n}\big{)}\sim\big{(}\hat{b}_{n}\big{)}$ everywhere in $\mathbb{R}_{++}^{m\times\infty}$, they are said to be universally indifferent.
3 A construction for combinatorial strategy
3.1 The benchmark strategy
In this section, a benchmark strategy, denoted as $\big{(}b_{n}^{*}\big{)}$, is formed using the concept in Definition 4 such that the preference $\big{(}b_{n}^{*}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ is more universal than the reverse one $\big{(}\ddot{b}_{n}\big{)}\succsim\big{(}b_{n}^{*}\big{)}$ in the sense of Claim 1. Although the strategy $\big{(}b_{n}^{*}\big{)}$ is unattainable as it relies on future data like the baseline $\big{(}\ddot{b}_{n}\big{)}$, we could create a combinatorial strategy $\big{(}b_{n}\big{)}$ using only past data such that the preference $\big{(}b_{n}\big{)}\sim\big{(}b_{n}^{*}\big{)}$ holds everywhere in $\mathbb{R}_{++}^{m\times\infty}$, making $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ more universal than the reverse preference by the transitivity of the preference relation.
Definition 4.
Given a set of component strategies $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$, for any constant combination $\lambda\coloneqq\left(\lambda_{1},...,\lambda_{k}\right)\in\mathcal{B}^{k}$, a corresponding constant combinatorial strategy $\big{(}b_{n}^{\lambda}\big{)}$ is formed. Subsequently, the best constant combination $\lambda_{n}\coloneqq\left(\lambda_{n,1},...,\lambda_{n,k}\right)$ at time $n$ is determined as follows:
$\lambda_{n}\coloneqq\operatorname*{argmax}_{\lambda\in\mathcal{B}^{k}}S_{n}\big{(}b_{n}^{\lambda}\big{)},\text{ where }b_{n}^{\lambda}\coloneqq{\displaystyle\sum_{\alpha=1}^{k}}\lambda_{\alpha}b_{n}^{\alpha},\,\forall n.$
Each constant combinatorial strategy allocates fixed proportions of capital into the corresponding component strategies. By utilizing this concept, we establish a benchmark strategy as the best merger in hindsight of all component strategies $\big{(}b_{n}^{\lambda}\big{)}$ similarly to Definition 3.
Definition 5.
(Benchmark strategy). The benchmark strategy, represented by $\big{(}b_{n}^{*}\big{)}$ henceforth, is defined as one satisfying $S_{n}\big{(}b_{n}^{*}\big{)}\coloneqq S_{n}\big{(}b_{n}^{\lambda_{n}}\big{)}$ for all $n$, which is simply the best-merged strategy of all component constant combinatorial strategies $\big{(}b_{n}^{\lambda}\big{)}$.
Definition 5 immediately results in $W_{n}\big{(}b_{n}^{*}\big{)}\geq W_{n}\big{(}\ddot{b}_{n}\big{)}$, so the preference $\big{(}b_{n}^{*}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ holds for any sequence $x_{1}^{\infty}$. If $\big{(}b_{n}^{*}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ for some sequences in $\mathbb{R}_{++}^{m\times\infty}$, any combinatorial strategy satisfying $\big{(}b_{n}\big{)}\sim\big{(}b_{n}^{*}\big{)}$ implies $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ due to the transitivity and asymmetry of the preference relation, thus meeting the Combination criterion. Therefore, it is sufficient to secure the preference $\big{(}b_{n}^{*}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ for a sequence $x_{1}^{\infty}$ by verifying if $\big{(}W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}\ddot{b}_{n}\big{)}\big{)}\nrightarrow 0$ as $n\to\infty$ due to $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}\land\big{(}b_{n}\big{)}\nsim\big{(}\ddot{b}_{n}\big{)}\Leftrightarrow\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$.
By definition, the constant combination mechanism exploits the inherent extent of difference in returns between the component strategies. If there is a significant discrepancy in the behaviors of capital growth among the component strategies, the maximizers $\lambda_{n}$ will not consistently remain close to the vertices as time develops. As a result, the benchmark strategy will infinitely often exceed the baseline strategy, and so $W_{n}\big{(}b_{n}^{*}\big{)}>W_{n}\big{(}\ddot{b}_{n}\big{)}$ at infinitely many $n$. The following proposition provides a basis for explaining the behavior of the benchmark strategy.
Proposition 3.
Given a set of strategies $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$, suppose there exist $h$ disjoint and non-empty sets indexed as $\mathcal{A}^{j}$, $\bigcup_{j\in\left[h\right]}\mathcal{A}^{j}=\left[k\right]$, and a set $\mathcal{H}\subseteq\left[k\right]$ such that $\left|\mathcal{H}\cap\mathcal{A}^{j}\right|=1$ for all $j\in\left[h\right]\coloneqq\big{\{}1,...,h\big{\}}$, which satisfy the following:
$\big{<}b_{n}^{\alpha^{j}},x_{n}\big{>}\geq\big{<}b_{n}^{\alpha},x_{n}\big{>},\text{ $\forall\alpha\in\mathcal{A}^{j}$,\,$\forall n$}\text{, where }\left\{\alpha^{j}\right\}\coloneqq\mathcal{H}\cap\mathcal{A}^{j},j\in\left[h\right].$
Let $\lambda\in\mathcal{B}^{k}$ and $\gamma\in\mathcal{B}^{h}$ denote the constant combinations of the component strategies in $\mathcal{A}$ and $\mathcal{H}$, respectively, then $S_{n}\big{(}b_{n}^{\lambda_{n}}\big{)}=S_{n}\big{(}b_{n}^{\gamma_{n}}\big{)}$ for all $n$.
Proof.
See Appendix. ∎
Proposition 3 states that if the set of component strategies can be separated into $h$ subsets, such that each subset contains a single strategy that never yields lower returns than the other strategies in the same subset, then the cumulative wealth of the baseline strategy will always be equal to that of the best merged one of only $h$ dominating strategies. Consequently, the maximizers $\lambda_{n}$ have only $h$ nonzero components regardless of the number of component strategies. Hence, if there exists a single strategy that generates higher returns than all others after a finite number of time periods, the maximizers $\lambda_{n}$ will remain close to a vertex of $\mathcal{B}^{k}$ as time develops, no matter how the remainders behave, so $\big{(}W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}\ddot{b}_{n}\big{)}\big{)}\rightarrow 0$.
3.2 Small scale of component strategies
In this section, we consider the case of a small number $k\geq 2$ for the set $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$ of component strategies. After introducing a specific benchmark strategy $\big{(}b_{n}^{*}\big{)}$ in the previous section, we propose now a combinatorial strategy $\big{(}b_{n}\big{)}$, which does not depend on future data, such that $\big{(}b_{n}\big{)}\sim\big{(}b_{n}^{*}\big{)}$ for all instances of sequences in $\mathbb{R}_{++}^{m\times\infty}$. This goal can be accomplished by demonstrating that the distance $\left|\log S_{n}\big{(}b_{n}^{*}\big{)}-\log S_{n}\big{(}b_{n}\big{)}\right|$ in the worst case of sequence $x_{1}^{n}$ is sublinear in the time variable $n$. To this end, we propose a combinatorial strategy as follows:
Strategy proposal: At each time $n$, the investor makes the following combination after observing the past cumulative wealths $S_{n-1}\big{(}b_{n-1}^{\alpha}\big{)}$ of $k$ component strategies:
$b_{1}\coloneqq{\displaystyle\dfrac{1}{k}{\displaystyle\sum_{i=1}^{k}b_{1}^{i}}}\text{ and }{\displaystyle b_{n}\coloneqq{\displaystyle\dfrac{{\displaystyle\int_{\mathcal{B}^{k}}{\displaystyle b_{n}}}S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}\mu\left(\lambda\right)\mathop{}\!\mathrm{d}\lambda}{{\displaystyle\int_{\mathcal{B}^{k}}S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}\mu\left(\lambda\right)\mathop{}\!\mathrm{d}\lambda}}},\,\forall n\geq 2,}$ (3.1)
where $\mu\left(\lambda\right)$ is the probability density with respect to the variable $\lambda\in\mathcal{B}^{k}$, and $\big{(}b_{n}^{\lambda}\big{)}$ is the strategy corresponding to a constant combination $\lambda$ of the component strategies.
If the density $\mu\left(\lambda\right)$ is uniform, the proposed combinatorial strategy given by (3.1) is always exceeded by the benchmark strategy due to:
[…]
$<S_{n}\big{(}b_{n}^{\lambda_{n}}\big{)}\int_{\mathcal{\mathcal{B}}^{k}}\mu\left(\lambda\right)\mathop{}\!\mathrm{d}\lambda=S_{n}\big{(}b_{n}^{*}\big{)},\,\forall n.$ (3.2)
In this case, the combinatorial strategy is equivalent to the exponentially weighted average forecaster with the learning rate parameter of one and the logarithmic loss function, as established in Cesa-Bianchi et al. (1997), Cesa-Bianchi and Lugosi (1999, 2006). Proposition 4, given below, establishes an upper bound on the positive difference $\log S_{n}\big{(}b_{n}^{*}\big{)}-\log S_{n}\big{(}b_{n}\big{)}$ in the worst case of the sequence $x_{1}^{n}$. It follows from the theorem for the minimax problem shown in Cover (1991), Cover and Ordentlich (1996), Cover and Thomas (2006) and Cesa-Bianchi and Lugosi (2006) [^2] Within the framework of Prediction with expert advice, as discussed in Cesa-Bianchi et al. (1997), Cesa-Bianchi and Lugosi (1999, 2006), the Universal portfolio is equivalent to the Laplace mixture forecaster, studied in Davisson (1973), Rissanen (1986), for the case of uniform distribution, and to the Krichevsky–Trofimov mixture forecaster, proposed by Krichevsky and Trofimov (1981), for the Dirichlet distribution in the Kelly betting market. It has also been investigated within another umbrella of Online portfolio selection, as in Ordentlich and Cover (1996), Helmbold et al. (1998), Blum and Kalai (1999), Fostera and Vohra (1999), Vovk and Watkins (1998), van Erven et al. (2020)..
Proposition 4.
If the combinatorial strategy $\big{(}b_{n}\big{)}$ is constructed according to (3.1) with the uniform probability density $\mu\left(\lambda\right)$ over the simplex $\mathcal{\mathcal{B}}^{k}$, then:
${\displaystyle\max_{x_{1}^{n}}\left(\log S_{n}\left(b_{n}^{*}\right)-\log S_{n}\left(b_{n}\right)\right)\leq\left(k-1\right)\log\left(n+1\right),\text{ $\forall n$}},$
where $\big{(}b_{n}^{*}\big{)}$ is the benchmark strategy of all component strategies in the set $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$.
Proof.
See Appendix. ∎
Due to Proposition 4 and the strict inequality (3.2), ${\displaystyle\big{(}{\textstyle W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}b_{n}\big{)}}\big{)}}\to 0$, i.e., $\big{(}b_{n}^{*}\big{)}\sim\big{(}b_{n}\big{)}$, for any sequence $x_{1}^{\infty}$ by the sandwich theorem. Note that within the framework of online portfolio optimization , there exist several other algorithms as surveyed in van Erven et al. (2020), that satisfy the sublinear upper bound for $\big{(}\log S_{n}\big{(}b_{n}^{*}\big{)}-\log S_{n}\big{(}b_{n}\big{)}\big{)}$. However, despite their computational efficiency, these algorithms rely on certain assumptions, which limit their guarantee in terms of bounding the minimax problem that is required to theoretically guarantee the preference $\big{(}b_{n}^{*}\big{)}\sim\big{(}b_{n}\big{)}$ for all sequences $x_{1}^{\infty}$. Moreover, our intentional choice of the construction for a combinatorial strategy according to (3.1) is to allow further flexibility in its algorithm. In the experiment section, we propose an accelerated variant for the combinatorial strategy using a time-variant density instead of the fixed uniform one. This acceleration method modifies the density by adaptively changing its support over time, enabling the combinatorial strategy to even exceed the cumulative wealth of the benchmark strategy.
3.3 Arbitrarily large scale of component strategies
When the scale of the set $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$ becomes larger, the computation of the proposed combinatorial strategy becomes increasingly difficult, especially at relatively large scales. Furthermore, the limit ${\displaystyle\big{(}{\textstyle W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}b_{n}\big{)}}\big{)}}\to 0$ is generally not guaranteed as $k\to\infty$, since the upper bound established in Proposition 4 is linear in $k$, meaning that $\left(k-1\right)\log\left(n+1\right)\to\infty$. Additionally, when $k\to\infty$, a convex combination of all the component portfolios results in atomic weights, rendering the combination meaningless. As a result, the combinatorial strategy reduces to nothing more than an exponentially weighted average of all component strategies, as formed by the formula (4.1) in the experiment section, thereby failing to meet the Combination criterion, as $\big{(}b_{n}^{*}\big{)}\sim\big{(}\ddot{b}_{n}\big{)}$ for any sequence $x_{1}^{\infty}$. To address this issue, we propose a modified combinatorial strategy below.
Strategy proposal: First of all, we form a finite number of base sets by partitioning an arbitrarily large set $\left[k\right]$ into $N$ smaller disjoint and nonempty sets, denoted as $\mathcal{A}^{j}$, $j\in\left[N\right]\coloneqq\left\{1,...,N\right\}$, $\bigcup_{j\in\left[N\right]}\mathcal{A}^{j}=\left[k\right]$. At each time period $n$, the investor decides the combination as follows:
$b_{1}\coloneqq{\displaystyle\dfrac{1}{N}{\displaystyle\sum_{i=1}^{N}\hat{b}_{1}^{i}}}\text{ and }b_{n}\coloneqq{\displaystyle\dfrac{{\displaystyle\int_{\mathcal{B}^{N}}{\displaystyle b_{n}}}S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}\mu\left(\lambda\right)\mathop{}\!\mathrm{d}\lambda}{{\displaystyle\int_{\mathcal{B}^{N}}S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}\mu\left(\lambda\right)\mathop{}\!\mathrm{d}\lambda}}},\,\forall n\geq 2,$ (3.3)
where $\mu\left(\lambda\right)$ is the probability density with respect to the variable $\lambda\in\mathcal{\mathcal{B}}^{N}$, and $\big{(}{\displaystyle b_{n}^{\lambda}\big{)}}$ denotes the strategy corresponding to a constant combination $\lambda$ of all $N$ representative strategies $\big{(}\hat{b}_{n}^{i}\big{)}$. In detail, each constant combinatorial portfolio $b_{n}^{\lambda}$ is constructed as follows:
$b_{n}^{\lambda}={\displaystyle\sum_{i=1}^{N}\lambda_{i}}\hat{b}_{n}^{i},\,\forall n,$
where the representative portfolios $\hat{b}_{n}^{i}$ are constructed as:
$\hat{b}_{1}^{i}\coloneqq{\displaystyle\sum_{\alpha\in\mathcal{A}^{i}}{\displaystyle b_{1}^{\alpha}}}\mu^{i}\big{(}\alpha\big{)},\,\forall i\in\left[N\right]\text{ and }\hat{b}_{n}^{i}\coloneqq\dfrac{{\displaystyle{\displaystyle\sum_{\alpha\in\mathcal{A}^{i}}}{\displaystyle b_{n}^{\alpha}}}S_{n-1}\left(b_{n-1}^{\alpha}\right)\mu^{i}\left(\alpha\right)}{{\displaystyle\sum_{\alpha\in\mathcal{A}^{i}}}{\displaystyle S_{n-1}\left(b_{n-1}^{\alpha}\right)\mu^{i}\left(\alpha\right)}},\forall n\geq 2,i\in\left[N\right],$ (3.4)
where each $\mu^{i}\left(\alpha\right)$ is the probability mass with respect to $\alpha\in\mathcal{A}^{i}$ and $i\in\left[N\right]$.
The proposed combinatorial strategy utilizes the partition to reduce dimensionality, where each base set represents a population of component strategies corresponding to a distribution. Particularly, the representative strategies $\big{(}\hat{b}_{n}^{i}\big{)}$, as per (3.4), represent aggregate behaviors of all the component strategies in each corresponding base set with respect to the mass $\mu^{i}\left(\alpha\right)$; then the combinatorial strategy $\big{(}b_{n}\big{)}$ is constructed based on these representative strategies according to (3.3), which is similar to the formula (3.1). In addition, concerning the allocation of wealth among the component strategies, let $P_{n}\big{(}b_{n}^{\alpha}\big{)}$ denote the time-dependent distribution of wealth over the corresponding component strategies $\big{(}b_{n}^{\alpha}\big{)}$; then, the capital allocated to each component portfolio $b_{n}^{\alpha}$ at time $n$ is given as follows:
$P_{n}\big{(}b_{n}^{\alpha}\big{)}={\displaystyle\dfrac{{\displaystyle\int_{\mathcal{B}^{k}}{\displaystyle\lambda_{i}}}S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}S_{n-1}\big{(}b_{n-1}^{\alpha}\big{)}\mu^{i}\big{(}\alpha\big{)}\mu\left(\lambda\right)\mathop{}\!\mathrm{d}\lambda}{{\displaystyle\int_{\mathcal{B}^{k}}\sum_{\beta\in\mathcal{A}^{i}}S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}{\displaystyle S_{n-1}\big{(}b_{n-1}^{\beta}\big{)}\mu^{i}\left(\beta\right)\mu\left(\lambda\right)}}\mathop{}\!\mathrm{d}\lambda}},\,\forall\alpha\in\mathcal{A}^{i},i\in\left[N\right].$
Thus, the return at time $n$ of the combinatorial portfolio $\big{(}b_{n}\big{)}$ is ${\displaystyle\sum}_{\alpha\in\left[k\right]}\left\langle b_{n}^{\alpha},x_{n}\right\rangle P_{n}\big{(}\alpha\big{)}$.
According to the construction of the proposed strategy, the benchmark strategy $\big{(}b_{n}^{*}\big{)}$ is no longer formed by the constant combinations of the component strategies but of the representative strategies of the base sets, so the worst-case upper bound derived in Proposition 4 does not apply. Instead, Proposition 5 derives the upper bound for the difference $\log S_{n}\big{(}\ddot{b}_{n}\big{)}-\log S_{n}\big{(}b_{n}\big{)}$ in the worst case of sequence $x_{1}^{n}$. However, contrary to Proposition 4, this worst-case difference is not always lower bounded by zero as in (3.2).
Proposition 5.
Consider a combinatorial strategy $\big{(}b_{n}\big{)}$ constructed according to (3.3) with the uniform density $\mu\left(\lambda\right)$ over the simplex $\mathcal{\mathcal{B}}^{N}$, and the minimal mass values ${\displaystyle\min_{i\in\left[N\right]}}\mu^{i}\big{(}\alpha_{n}^{i}\big{)}\eqqcolon\epsilon_{n}>0$ at $\alpha_{n}^{i}$, which are defined in hindsight as:
$\alpha_{n}^{i}=\operatorname*{argmax}_{\alpha\in\mathcal{A}^{i}}S_{n}\big{(}b_{n}^{\alpha}\big{)},\,\forall i\in\left[N\right],\,n\geq 2,$
and the probability mass $\mu^{i}\left(\alpha\right)$ for the representative portfolio $\hat{b}_{n}^{i}$ according to (3.4). We have:
${\displaystyle\max_{x_{1}^{n}}\big{(}\log S_{n}\big{(}\ddot{b}_{n}\big{)}-\log S_{n}\big{(}b_{n}\big{)}\big{)}\leq\left(N-1\right)\log\left(n+1\right)-\log\epsilon_{n}},\text{ $\forall n$},$
where $\big{(}\ddot{b}_{n}\big{)}$ is the baseline strategy of all component strategies in the set $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$.
Proof.
See Appendix. ∎
If the minimal mass values $\epsilon_{n}$ are chosen such that $n^{-1}\log\epsilon_{n}\to 0$, then by Proposition 5, we have $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$ for any sequence $x_{1}^{\infty}$, as the upper bound converges to zero. The case of representative strategies with uniform probability mass $\mu^{i}\left(\alpha\right)$ over each countable base set is an instance ensuring such convergence. As we noted above, the infimum limit is not bounded by zero, so there exist sequences $x_{1}^{\infty}$ where the combinatorial strategy $\big{(}b_{n}\big{)}$ is not always exceeded by the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$. The remark below specifies conditions to guarantee the Combination criterion. In addition, the time required for the combinatorial strategy to exceed all component strategies increases significantly with larger numbers of base sets and their respective scales. Additionally, the partitioning into base sets affects the condition $\big{(}W_{n}\big{(}b_{n}\big{)}-W_{n}\big{(}\dot{b}_{t}\big{)}\big{)}\nrightarrow 0$ and the evolutions of representative strategies’ cumulative wealths. Furthermore, the combinatorial strategy reduces to the one proposed for a small number $k$ of component strategies by equating $N=k$.
Remark.
Let $\big{(}\dot{b}_{n}\big{)}$ denote the best merged strategy of all the representative strategies, as similarly defined as the baseline strategy in Definition 3. The benchmark strategy $\big{(}b_{n}^{*}\big{)}$ is defined as the constant combination of the representative strategies, not the component strategies. If $n^{-1}\log\epsilon_{n}\to 0$, then $\big{(}\ddot{b}_{n}\big{)}\sim\big{(}\dot{b}_{n}\big{)}$ for any sequence $x_{1}^{\infty}$; moreover, for sequences such that $\big{(}W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}\dot{b}_{n}\big{)}\big{)}\nrightarrow 0$, then $\big{(}b_{n}\big{)}\succ\big{(}\dot{b}_{n}\big{)}$ so $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ which satisfies the Combination criterion, due to the transitivity and asymmetry of the preference relation.
4 Numerical experiments
The integrals of the combinatorial portfolios corresponding to (3.3) or (3.1) and (3.4) can be numerically approximated by Riemann sums that partition the simplices $\mathcal{B}^{N}$ and $\mathcal{B}^{k}$ into finitely discrete grids. Essentially, a finer partition can enhance the approximation, but it comes at the cost of increasing computational complexity, especially with high-dimensional simplices. Figure 2 provides an example from the experiment section, showcasing the numerical approximation of the investor’s combinatorial strategy using single-step size discretization for the 1-dimensional simplex. It illustrates how more discretization points and smaller step sizes can improve cumulative wealth. Additionally, various step sizes and discretization points can be combined for further refinement. It is important to note that while the theoretically optimal constant combinations according to Definition 4 may not align with discretization points, fine partitions can ensure that numerical maximizers closely approximate theoretical values. However, enhancing numerical approximation is not the primary focus of this experimental section.
Figure 2: Final cumulative wealth plotted against discretization points and step sizes for approximating the combinatorial strategy with two component strategies.
4.1 Data and testing strategies
We experiment with six diverse stocks over a 27-year span, from December 31, 1992, to December 31, 2019, covering $6798$ trading days. This period encapsulates major market events like the dot-com bubble, the 2008 financial crisis, and pre-COVID-19 pandemic, as illustrated in Figure 3. The selected stocks are blue-chip symbols, ranging across several sectors like industrials, energy, technology, and consumer defensive, traded on the NYSE and NASDAQ exchanges. They include Advanced Micro Devices, Inc. (AMD), The Boeing Company (BA), Honeywell International Inc. (HON), The Coca-Cola Company (KO), 3M Company (MMM), and Exxon Mobil Corporation (XOM). Their varied behaviors during historic events challenge prediction, making them ideal for testing our strategies. We used daily adjusted closing prices in USD provided by The Center for Research in Security Prices, LLC (CRSP).
Figure 3: Adjusted closing prices in USD of the stocks from December 31, 1992, to December 31, 2019.
Market scenario. Let’s consider a market where many fund managers have different risk profiles for their investment strategies, and an investor can observe them as the components of a combinatorial strategy. We consider the sequences of Mean-Variance portfolios corresponding to the managers’ risk aversion levels $\alpha\in\mathbb{R}_{+}$, which are defined as follows, as component strategies:
$b_{n}^{\alpha}\coloneqq\operatorname*{argmax}_{b\in\mathcal{B}^{m}}\big{(}\alpha\left\langle b,\hat{\mu}_{n}\right\rangle-\big{<}b,\hat{\varSigma}_{n}b\big{>}\big{)},\forall n,$
where $\hat{\mu}_{n}$ and $\hat{\varSigma}_{n}$ denote the estimators for the expected values and covariance matrix of the assets’ returns at time $n$. Specifically, we further assume that $\hat{\mu}_{n}$ and ${\displaystyle\hat{\varSigma}_{n}}$ are estimated using a mean-reversion approach that utilizes past prices of six assets within a moving window as:
$\hat{\mu}_{n}$ $\coloneqq\frac{{\displaystyle 1}}{J}\stackrel{{\scriptstyle[}}{{i}}=n-J]{n-1}{\sum}x_{i}\text{ \,\,and\,\, }\hat{\varSigma}_{n}\left(i,j\right)\coloneqq{\displaystyle\frac{1}{J-1}}\stackrel{{\scriptstyle[}}{{h}}=n-J]{n-1}{\sum}\big{(}x_{h,i}-\hat{\mu}_{n,i}\big{)}\big{(}x_{h,j}-\hat{\mu}_{n,j}\big{)},\text{ $\forall i,j$$\in\left[6\right],\forall n>J,$}$
where $J$ is the scrolling window length set to four weeks $\left(J=20\right)$ for the entire investment simulation, starting from January 29, 1993, which is considered as the origin day of time $n=1$. The M-V portfolios are estimated numerically using the OSQP solver by Stellato et al. (2020), implemented within the CVXPY module for Python by Diamond and Boyd (2016) and Agrawal et al. (2018).
Testing strategies. We propose further an accelerated variant of the combinatorial strategy and assess their performances against various other (online learning) multi-strategies. The involved strategies and their corresponding abbreviations are summarized as follows:
  • $\bullet$
    The proposed combinatorial strategy according to (3.3) or (3.1), and (3.4) with uniform $\mu\left(\cdot\right)$ and $\mu^{i}\left(\cdot\right)$. We name it as Universally Combinatorial strategy (UC) with associated notation $\big{(}b_{n}^{\text{UC}}\big{)}$, inspired by the Universal Portfolio proposed by Cover (1991).
  • $\bullet$
    The aforementioned Mean-Variance strategies (M-V), denoted as $\big{(}b_{n}^{\alpha}\big{)}$ with various coefficients $\alpha\in\mathbb{R}_{+}$, which serve as the component strategies for all the tested multi-strategies.
  • $\bullet$
    The Weighted Average Exponential strategy (WAE). In detail, the WAE portfolios, denoted by $\big{(}b_{n}^{\text{WAE}}\big{)}$, combine $k$ component portfolios in the set $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$, as follows:
    $b_{1}^{\text{WAE}}=\dfrac{1}{k}\sum_{\alpha\in\left[k\right]}b_{1}^{\alpha}\text{ and }b_{n}^{\text{WAE}}={\displaystyle{\displaystyle\dfrac{{\displaystyle\mathop{\sum_{\alpha\in\left[k\right]}b_{n}^{\alpha}}S_{n-1}\big{(}b_{n-1}^{\alpha}\big{)}}}{{\displaystyle\mathop{\sum_{\alpha\in\left[k\right]}}S_{n-1}\big{(}b_{n-1}^{\alpha}\big{)}}},\text{ $\forall n\geq 2$}.}}$ (4.1)
  • $\bullet$
    The Following-the-Leader strategy (FL), denoted by $b_{n}^{\text{FL}}$, uses the best component strategy, among $k$ component strategies in $\big{\{}\big{(}b_{n}^{\alpha}\big{)}:\alpha\in\left[k\right]\big{\}}$, in hindsight after $n-1$ periods for the next period $n$. Specifically, it is defined as follows:
    $b_{1}^{\text{FL}}=\dfrac{1}{k}\sum_{\alpha\in\left[k\right]}b_{1}^{\alpha}\text{ \,and\, }b_{n}^{\text{FL}}=b_{n-1}^{\alpha_{n-1}},\text{ where }\text{$\alpha_{n-1}$}\coloneqq\operatorname*{argmax}_{\alpha\in\left[k\right]}S_{n-1}\big{(}b_{n-1}^{\alpha}\big{)},\forall n\geq 2.$
  • $\bullet$
    The accelerated variant of the combinatorial strategy uses a time-varying density in place of the time-invariant uniform one over the simplex $\mathcal{B}^{k}$. Namely, we introduce the Universal combination on the Winners (UC-W) as a UC strategy, where the uniform density $\mu_{n}$ has support on $B_{n}\subset\mathcal{B}^{k}$ at period $n\geq 2$, defined as follows:
    $B_{n}\coloneqq\left\{\lambda\in\mathcal{B}^{k}:\,S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}>S_{n-1}\big{(}b_{n-1}^{\bar{\lambda}}\big{)},\,\forall\bar{\lambda}\in\mathcal{B}^{k}\backslash B_{n}\right\},\forall n\geq 2.$ (4.2)
    Accordingly, the Universal combination on the Losers (UC-L) is formed using the subsets:
    $\bar{B}_{n}\coloneqq\left\{\lambda\in\mathcal{B}^{k}:\,S_{n-1}\big{(}b_{n-1}^{\lambda}\big{)}<S_{n-1}\big{(}b_{n-1}^{\bar{\lambda}}\big{)},\,\forall\bar{\lambda}\in\mathcal{B}^{k}\backslash\bar{B}_{n}\right\},\forall n\geq 2.$ (4.3)
    This method is applicable to the simplex $\mathcal{B}^{N}$ for the strategy constructed according to (3.3). The UC-W and UC-L strategies are also computed using Riemann approximation.
The experiments illustrate the evolution of cumulative wealths of the testing strategies through several graphics. Specifically, we investigate whether they exhibit a strict preference over the baseline strategies by eventually exceeding all the M-V component strategies. Additionally, we report common performance measures for the testing strategies, including their final cumulative wealth, growth rate, average return (i.e., empirical expected return), and empirical Sharpe ratio.
4.2 Experiments with small scales of component strategies
In the first four experiments, we implement different small numbers of component M-V strategies to examine the impact of scales on the behaviors of the multi-strategies, including $\left\{0.005,1\right\}$, $\left\{0.002,0.02,100\right\}$, $\left\{0.05,0.2,0.5,0.8,1\right\}$ and $\left\{0.002,0.05\right\}$ for risk aversions in Experiments 1, 2, 3, and 4, respectively. Generally, Figure 4 illustrates the UC strategies eventually exceed all the component M-V strategies, except in Experiment 4. The strategies WAE and FL are always exceeded by the best M-V strategies in all cases. The results indicate that the FL strategies demonstrate relative effectiveness if there exists a single M-V strategy that almost always exceeds the others after $N$ periods, then $\big{(}\ddot{b}_{n}\big{)}\sim\big{(}b_{n}^{\text{FL}}\big{)}$ due to the following:
$\lim_{n\to\infty}\big{(}W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}^{\text{FL}}\big{)}\big{)}=\lim_{n\to\infty}\frac{1}{n}\big{(}\log S_{N+1}\big{(}\ddot{b}_{N+1}\big{)}-\log S_{N+1}\big{(}b_{N+1}^{\text{FL}}\big{)}\big{)}=0.\vspace{0.25ex}$
However, the mechanisms of the FL and WAE strategies inherently constrain their growth rates, being bounded by the baseline strategy. Consequently, despite the possibility of an indifferent preference with the latter, the cumulative wealth evolutions of the two former strategies are worse. These experiments emphasize that attaining the same asymptotic growth rate as the best component strategy does not guarantee a multi-strategy not being the worst performer.
Table 1 presents performance metrics for the strategies involved, while Figure 5 displays the final cumulative wealth against the numerical constant combination. The Sharpe ratios of the multi-strategies do not surpass the best among the component M-V strategies. However, the UC strategies offer the most benefit as they achieve the highest final cumulative wealths in the first three experiments, with relatively small tradeoffs in terms of the Sharpe ratios. In contrast, all M-V strategies that yield higher final wealth and mean of return experience significant decreases in Sharpe ratios. These favorable performances of the UC strategies are attained through approximations with respect to the numerical combinations $\lambda$. Although their performances could potentially improve with a finer grid, Figure 5 demonstrates that such marginal improvement is not worthwhile, as the computation may become infeasible for a large scale of component strategies.
Figure 4: Evolutions of the multi-strategies over time in Experiments 1, 2, 3, and 4 (from left-to-right and top-to-bottom order).
Table 1: Performance measures for the strategies in Experiments 1, 2, 3 and 4.
Strategy Final wealth Average growth rate Average return Sharpe ratio
M-V ($\alpha=0.002$) 21.081643 0.000450 1.000511 90.884103
M-V ($\alpha=0.005$) 19.626578 0.000439 1.000500 90.777017
M-V ($\alpha=0.02$) 15.270352 0.000402 1.000468 87.209390
M-V ($\alpha=0.05$) 14.951959 0.000399 1.000482 77.889510
M-V ($\alpha=0.2$) 20.864838 0.000448 1.000615 54.677952
M-V ($\alpha=0.5$) 19.022646 0.000435 1.000700 43.451483
M-V ($\alpha=0.8$) 18.728020 0.000432 1.000737 40.520186
M-V ($\alpha=1$) 16.503138 0.000414 1.000734 39.543545
M-V ($\alpha=100$) 26.641673 0.000484 1.000849 37.017071
Experiment 1 with risk aversion $\alpha\in\left\{0.005,1\right\}$
UC 24.178108 0.000470 1.000580 67.419175
WAE 18.060145 0.000427 1.000513 76.449559
FL 16.015280 0.000409 1.000474 88.246262
Experiment 2 with risk aversion $\alpha\in\left\{0.002,0.02,100\right\}$
UC 29.192687 0.000498 1.000597 71.217927
WAE 20.982959 0.000449 1.000535 76.247127
FL 8.8740965 0.000322 1.000416 73.280794
Experiment 3 with risk aversion $\alpha\in\left\{0.05,0.2,0.5,0.8,1\right\}$
UC 20.887842 0.000448 1.000645 50.409559
WAE 18.014120 0.000427 1.000614 51.686007
FL 4.3021720 0.000215 1.000340 63.518664
Experiment 4 with risk aversion $\alpha\in\left\{0.002,0.05\right\}$
UC 18.401030 0.000430 1.000493 88.740964
WAE 18.016801 0.000427 1.000489 89.929546
FL 20.401686 0.000445 1.000506 90.848864
Note. The UC strategies are approximated by 2001 discretization points by step 0.0005 in Experiments 1 and 4, 4947 discretization points by step 0.01 in Experiment 2, and 9821 discretization points by step 0.05 in Experiment 3.
The first row of Figure 6 illustrates the growth rate differences between the numerically computed benchmark strategy $\big{(}b_{n}^{*}\big{)}$ and the UC strategy $\big{(}b_{n}^{\text{UC}}\big{)}$, as well as between the benchmark strategy and the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$ of all related M-V strategies. The Combination criterion is satisfied in the first three experiments but not in the remainder, as $\big{(}W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}^{*}\big{)}\big{)}\to 0$ as $n$ increases. In Experiment 1, the M-V strategy $\big{(}b_{n}^{0.005}\big{)}$ almost always exceeds the M-V strategy $\big{(}b_{n}^{1}\big{)}$, but the UC strategy is able to exceed the former. Conversely, in Experiment 4, the UC strategy never exceeds the best M-V strategy $\big{(}b_{n}^{0.002}\big{)}$, which almost always exceeds the M-V strategy $\big{(}b_{n}^{0.05}\big{)}$. In the two remaining experiments, the UC strategy is capable of exceeding the best M-V strategies, where one component strategy experiences strong fluctuations while the other two experience stable growth over time in Experiment 2, and five comparable component strategies continuously compete for the leading positions over time in Experiment 4.
Refer to caption Refer to captionRefer to captionRefer to caption
  Experiment 1    Experiment 2
Refer to caption Refer to captionRefer to captionRefer to caption
  Experiment 4    Experiment 3
Figure 5: The constant combination $\lambda$ plotted against the final cumulative wealth with respect to the discretization. The best constant combinations in Experiment 1, 2, 3, and 4 are $\left(0.55,0.45\right)$, $\left(0.44,0,0.56\right)$, $\left(0.3,0.2,0,0.5,0\right)$ and $\left(1,0\right)$, respectively.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
   Experiment 1    Experiment 2    Experiment 3    Experiment 4
Figure 6: Upper row: differences over time between the growth rates of the strategies $\big{(}\ddot{b}_{n}\big{)}$, $\big{(}b_{n}^{*}\big{)}$ and $\big{(}b_{n}^{\text{UC}}\big{)}$. Lower row: upper bound over time according to Proposition 4.
The second row of Figure 6 depicts the upper bound for the difference between the two numerically computed strategies, $\log S_{n}\big{(}b_{n}^{*}\big{)}-\log S_{n}\big{(}b_{n}^{\text{UC}}\big{)}$, as established in Proposition 4. Furthermore, in order to fully comprehend the behavior of the benchmark strategy $\big{(}b_{n}^{*}\big{)}$, the first row of Figure 7 exhibits the best numerically constant combinations over time with respect to different discretization step sizes in each experiment. The first three graphics indicate the presence of only two non-zero components of the time-variant maximizer $\lambda_{n}$, while the last graphic shows only one non-zero component. This is due to the M-V strategy $\big{(}b_{n}^{0.002}\big{)}$ consistently dominating the M-V strategy $\big{(}b_{n}^{0.05}\big{)}$ in the sense of Proposition 3, which explains the violation of the Combination criterion in Experiment 4. In contrast, the second row of Figure 7 illustrates the proportions in the UC strategy that the investor allocates to each component M-V strategy over time, showing that all of them are consistently invested in throughout.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
   Experiment 1    Experiment 2    Experiment 3    Experiment 4
Figure 7: Upper row: the best constant combinations $\lambda_{n}$ of the component M-V strategies over time with respect to the discretization. Lower row: allocations of capital proportions into the component strategies, with $P_{t}$ representing the allocations over time $t$.
Universally combinatorial strategy with acceleration. The results of the first four experiments emphasize the necessity of accelerating the UC strategy to enhance the converging speed of its growth rate to that of the benchmark strategy or even surpassing it. Since the UC strategy allocates capital to all component strategies over time, it includes unnecessarily those with inferior performance. Particularly, the underperformance of the UC strategy compared to the significantly better FL strategy in Experiment 4 justifies our motivation to incorporate the mechanism of the latter into the former. To address this issue, the proposed UC-L and UC-W strategies aim to reduce the amount of capital invested in underperforming component strategies. However, unlike the original UC strategy, these accelerated strategies do not come with a theoretical guarantee. Therefore, the UC strategy will be tested alongside both of its variants UC-L and UC-W in Experiments 5, 6, 7, and 8 to empirically examine their performances. In accordance with the definitions of UC-L and UC-W strategies, we consider the time-variant sets $B_{n}$ and $\bar{B}_{n}$ of constant combinations, which are defined in (4.2) and (4.3), in the top and bottom $30\%$ and $50\%$ of cumulative wealth from the previous time period $n-1$ among all combinations in the simplex $\mathcal{B}^{k}$. Only in Experiment 8, we further consider an additional UC-W strategy with the top $5\%$ of combinations. The performances of these strategies are presented in Figure 8 and Table 2.
Figure 8: Evolutions of the accelerated combinatorial strategies over time in Experiments 5, 6, 7, and 8 (from left-to-right and top-to-bottom order).
Table 2: Performance measures for the strategies in Experiments 5, 6, 7 and 8.
Strategy Final wealth Average growth rate Average return Sharpe ratio
Experiment 5 with risk aversion $\alpha\in\left\{0.005,1\right\}$
UC 24.178108 0.000470 1.000580 67.419175
UC-L (30%) 24.414120 0.000471 1.000710 45.790848
UC-L (50%) 25.333606 0.000477 1.000668 51.143733
UC-W (30%) 20.393190 0.000445 1.000519 82.463092
UC-W (50%) 23.260910 0.000464 1.000545 78.922235
Experiment 6 with risk aversion $\alpha\in\left\{0.002,0.02,100\right\}$
UC 29.192687 0.000498 1.000597 71.217927
UC-L (30%) 38.484047 0.000539 1.000668 62.238247
UC-L (50%) 36.758159 0.000532 1.000646 66.163429
UC-W (30%) 23.426020 0.000465 1.000559 72.989659
UC-W (50%) 24.483100 0.000472 1.000567 72.393094
Experiment 7 with risk aversion $\alpha\in\left\{0.05,0.2,0.5,0.8,1\right\}$
UC 20.887842 0.000448 1.000645 50.409559
UC-L (30%) 30.182974 0.000503 1.000741 45.813842
UC-L (50%) 26.442502 0.000483 1.000712 46.693717
UC-W (30%) 15.468676 0.000404 1.000564 55.928931
UC-W (50%) 16.903366 0.000417 1.000590 53.769018
Experiment 8 with risk aversion $\alpha\in\left\{0.002,0.05\right\}$
UC 18.401030 0.000430 1.000493 88.740964
UC-L (30%) 16.480301 0.000413 1.000489 81.431545
UC-L (50%) 17.113231 0.000419 1.000490 83.761159
UC-W (5%) 20.299331 0.000444 1.000505 90.939658
UC-W (30%) 20.010540 0.000442 1.000503 91.026628
UC-W (50%) 19.641169 0.000439 1.000500 90.694303
Note. The UC strategies are approximated by 2001 discretization points by step 0.0005 in Experiments 5 and 8, 4947 discretization points by step 0.01 in Experiment 6, and 9821 discretization points by step 0.05 in Experiment 7.
The results presented in Figure 8 align with our expectations. Since the FL strategy works in Experiment 4, it results in the UC-W ($5\%$) strategy demonstrating the most favorable evolution of cumulative wealth compared to the other tested strategies in the same scenario in Experiment 8. This can be attributed to the almost always exceedance of the M-V strategy $\big{(}b_{n}^{0.002}\big{)}$ over the M-V strategy $\big{(}b_{n}^{0.05}\big{)}$, which causes a combinatorial strategy corresponding to a constant combination closer to the best one, $\lambda_{n}=(1,0)$ for all $n$, to yield higher cumulative wealth. Consequently, the UC-W strategy performs better with a small subset $B_{n}$ that includes the point $(1,0)$, while the UC-L strategy with a small subset $\bar{B}_{n}$ that includes the point $(0,1)$ leads to a decline in cumulative wealth. However, in the other three experiments, the UC-W strategies underperform, whereas the UC-L strategies show significant improvement. A possible explanation for this behavior is similar to the argument explaining the poor performance of the FL strategy in the corresponding scenarios. When strongly fluctuating strategies are present among the component M-V strategies, relying solely on constant combination strategies that previously led in terms of cumulative wealth becomes less reliable. Additionally, Table 2 reaffirms a characteristic inherent in the UC-W and UC-L strategies derived from the UC strategy, namely, that a strategy with higher final cumulative wealth tends to produce a slight decrease in its Sharpe ratio.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
 Experiment 5: UC-L (30%)  Experiment 6: UC-L (30%)  Experiment 7: UC-L (30%)   Experiment 8: UC-W (5%)
Figure 9: Upper row: allocations of capital proportions into the component M-V strategies, with $P_{t}$ representing the allocations over time $t$. Lower row: differences over time between the growth rates of the strategies $\big{(}\ddot{b}_{n}\big{)}$, $\big{(}b_{n}^{*}\big{)}$ and $\big{(}b_{n}^{\text{UC-L}}\big{)}$ or $\big{(}b_{n}^{\text{UC-W}}\big{)}$.
The acceleration results in a noticeable change in the capital allocation, as depicted in the first row of 9. In Experiment 8, the UC-W ($5\%$) strategy primarily allocates capital to the M-V strategy $\big{(}b_{n}^{0.002}\big{)}$, which closely aligns with the optimal constant combination $(1,0)$ over time. Conversely, in the remaining three experiments, the UC-L ($30\%$) strategy exhibits significant fluctuations in the allocated proportions among the involved component M-V strategies over time, in contrast to the gradual changes observed in the proportions of the UC strategy. These substantial fluctuations in allocated proportions subsequently impact the differences in growth rate between the benchmark strategy and the accelerated strategies, as shown in the second row of Figure 9. The first three graphs indicate that the growth rate $W_{n}\big{(}b_{n}^{\text{UC-L}}\big{)}$ is not bounded by $W_{n}\big{(}b_{n}^{*}\big{)}$, whereas $W_{n}\big{(}b_{n}^{\text{UC-W}}\big{)}$ approaches the upper bound $W_{n}\big{(}b_{n}^{*}\big{)}$ in the final graph. The accelerated UC-L strategy in Experiments 6 and 7 quickly surpasses the benchmark strategy, likely due to the number of component M-V strategies that exhibit significant fluctuations in cumulative wealth evolution. In Experiment 5, the consistent exceeding of the M-V strategy $\big{(}b_{n}^{0.005}\big{)}$ over the M-V strategy $\big{(}b_{n}^{1}\big{)}$ enables the UC-W strategies, which allocate their capital to small sets of constant combination strategies that previously led in cumulative wealth, to perform better than the UC-L strategies most of the time. However, as the M-V strategy $\big{(}b_{n}^{1}\big{)}$ experiences stronger fluctuations after about $5000$ time periods, the UC-L strategies gain momentum and eventually surpass the UC-W strategies. In Experiments 6 and 7, characterized by significant fluctuations in at least one M-V strategy and frequent changes in the leading position, the UC-L strategy, which consistently allocates its capital to a small set of constant combination strategies that do not lead in cumulative wealth but provide stability in returns, quickly surpasses the benchmark strategy.
4.3 Experiments with large scales of component strategies
After conducting eight experiments in the previous section, we extend our investigation to the UC strategy given by (3.3) and its accelerated variant in a scenario involving a large scale of component M-V strategies. For this analysis, we consider a set $\mathcal{A}$ of risk aversion coefficients encompassing values ranging from $0.001$ to $0.009$ with a step size of $0.001$, from $0.01$ to $0.09$ with a step size of $0.01$, from $0.1$ to $0.9$ with a step size of $0.1$, and from $1$ to $10$ with a step size of $0.5$. In Experiment 9, we examine the impact of partitioning $\mathcal{A}$ into multiple base sets and their respective sizes on the performance and behavior of the UC strategy. To investigate this further, we analyze partitions into two, four, and six base sets. Specifically, in the case of two base sets, our analysis includes examining balanced sets with equal sizes spanning from $0.001$ to $0.5$ and from $0.6$ to $10$, as well as unbalanced sets covering the ranges of $0.001$ to $0.02$ and $0.03$ to $10$. In the case of four base sets, we divide the coefficients from $0.001$ to $5$ into three equal-sized sets, while assigning the remainder to the fourth set. In a similar vein, for six base sets, we evenly segment the coefficients ranging from $0.001$ to $7$ into the first five sets, while leaving the remainder for the final set. Following Experiment 9, we proceed to implement accelerated strategies tailored to the scenario involving six base sets in Experiment 10.
Figure 10: Left: evolutions of the combinatorial strategies over time with different base sets in Experiment 9. Right: evolutions of the accelerated combinatorial strategies over time with six base sets in Experiment 10. The blurred background depicts the evolutions of the component M-V strategies.
Table 3: Performance measures for the strategies in Experiments 9 and 10.
Strategy Final wealth Average growth rate Average return Sharpe ratio
M-V ($\alpha=0.001$) 21.753985 0.000454 1.000515 90.837290
M-V ($\alpha=0.06$) 13.955013 0.000389 1.000477 75.186736
Experiment 9 with risk aversion $\alpha\in\mathcal{A}$
UC (2 unbalanced base sets) 21.991619 0.000456 1.000552 72.286540
UC (2 base sets) 22.882642 0.000462 1.000604 59.289323
UC (4 base sets) 23.847455 0.000468 1.000627 56.137547
UC (6 base sets) 24.360797 0.000471 1.000633 55.534858
WAE 17.919960 0.000426 1.000551 63.130386
Experiment 10 with risk aversion $\alpha\in\mathcal{A}$ and 6 base sets partition
UC-L (30%) 29.405695 0.000499 1.000725 46.965682
UC-L (50%) 27.598225 0.000490 1.000697 49.072607
UC-W (30%) 21.596212 0.000453 1.000578 63.439990
UC-W (50%) 22.317208 0.000458 1.000594 60.699273
Note. The UC, UC-L, and UC-W strategies are approximated by 2001 discretization points by step 0.0005, 8605 discretization points by step 0.0277 and 14831 discretization points by step 0.0666 for the partitions into 2, 4 and 6 base sets, correspondingly. Two M-V strategies corresponding to the risk aversion levels 0.001 and 0.06 are the best and the worst ones in terms of the final cumulative wealth.
Performance measures of the UC strategy, along with a WAE strategy applied to all component M-V strategies, are presented in Figure 10 and Table 3. The results demonstrate that the WAE strategy performs poorly in comparison to the UC strategy across all partitions, despite its higher Sharpe ratio. This is because the growth rate of the WAE strategy is bounded by that of the best component M-V strategy. Overall, all partitions lead to the outperformance of the UC strategy over the best M-V strategy in terms of cumulative wealth. However, the partition into unbalanced base sets shows relatively lower cumulative wealth compared to more evenly segmented base sets. Furthermore, increasing the number of base sets while reducing the size of each set simultaneously improves both the average growth rate, implying higher final cumulative wealth, and the average return, but results in a lower Sharpe ratio. These findings align with the characteristics observed in the previous eight experiments of the UC strategy. In more detail, regarding cumulative wealth, the UC strategy with two unbalanced base sets merely matches the cumulative wealth of the best M-V strategy after approximately $6600$ time periods, while the UC strategy with two balanced base sets eventually exceeds it. After the same amount of time, the UC strategies with four and six base sets exceed the best M-V strategy, with slightly more favorable cumulative wealth for the latter partitioning. This distinction is clearly depicted in Figure 11.
In the first row of Figure 11, it is evident that a higher number of base sets leads to a slower convergence speed for the growth rates of the benchmark and UC strategies. Conversely, partitioning into smaller-sized base sets can rapidly lift the difference $W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}\ddot{b}_{n}\big{)}$ above zero over time, enabling the UC strategy to exceed the baseline strategy faster. However, this difference may be impeded from exceeding zero by partitioning into strongly unbalanced base sets, as observed in the first upper graph. This phenomenon is caused by the behaviors of the representative strategies of the respective base sets, as a base set may contain numerous M-V strategies with markedly different behaviors from the others in the same set. Additionally, the second row of Figure 11 illustrates the upper bound for the difference $W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}^{\text{UC}}\big{)}$, established in Proposition 5, with respect to the numerically approximated UC strategy. It shows that, in general, the average difference over time converges to zero more slowly when the number of base sets is higher. The results of Experiment 9 emphasize the importance of appropriately tuning the partition for the set $\mathcal{A}$, such that the number of base sets should not be large, while each base set should mainly contain component strategies that behave in a similar manner.
Refer to caption Refer to caption Refer to caption Refer to caption
 Refer to caption  Refer to caption  Refer to caption  Refer to caption
   2 unbalanced base sets      2 base sets      4 base sets      6 base sets
Figure 11: Upper row: differences over time between the growth rates of the strategies $\big{(}\ddot{b}_{n}\big{)}$, $\big{(}b_{n}^{*}\big{)}$ and $\big{(}b_{n}^{\text{UC}}\big{)}$. Lower row: upper bound over time according to Proposition 5.
Refer to caption Refer to caption Refer to caption Refer to caption
    UC: 2 base sets     UC: 4 base sets     UC: 6 base sets    UC-L (30%): 6 base sets
Figure 12: Allocations of capital proportions into the component M-V strategies over time (from bottom-to-top: risk aversion levels from $0.001$ to $10$).
In Experiment 10, we conduct the UC-L and UC-W strategies alongside the UC strategy using a partition of six base sets. The computation procedure for the UC-L and UC-W strategies remains unchanged, except for the replacement of component M-V strategies with representative strategies. The results of this experiment are presented in Figure 11 and Table 3. Similar to the case with a small scale of component strategies, the UC-L strategy with a small set $\bar{B}_{n}$ exhibits higher cumulative wealth and average return compared to the other two, albeit with a slight decrease in the Sharpe ratio. Figure 12 illustrates the allocation proportions of capital to each component M-V strategy over time by the UC-L ($30\%$) strategy, in comparison to the original UC strategy with all partitions except the unbalanced one. The graphs show that the proportions experience lower variability as the number of base sets increases, while demonstrating stronger fluctuations under acceleration for the UC strategy. This phenomenon can be explained using the results from the experiments conducted with small scales of risk aversion levels. Specifically, with fewer base sets, a larger number of component strategies are adjusted relative to the weights assigned to their corresponding representative strategies, and vice versa. Moreover, changes in the weights of each representative strategy are more gradual when their number is higher, as depicted in the second row of Figure 7. Therefore, the extent of fluctuation in the proportions of the UC strategy decreases as the number of base sets increases. However, since acceleration causes strong variations in the weights of representative strategies, the proportions of the component strategies also exhibit significant fluctuations, as shown in the last graph for the UC-L strategy.
5 Summary and concluding Remarks
In this paper, we consider an investor in the market who prioritizes wealth accumulation and possesses the ability to learn several observable strategies represented as sequences of no-short portfolios, which are then ensembled into a multi-strategy. We establish a distribution-free preference framework based on asymptotic growth rate to ensure that the multi-strategy eventually exceeds all component strategies simultaneously in terms of cumulative wealth, provided it is strictly preferred to the baseline strategy as a merger of all components. Moreover, given the unknown possibilities of the infinite sequence of market assets’ returns, the combinatorial strategy must exhibit weak preference everywhere with strict preference at some instances of the sequence compared to the baseline strategy. Subsequently, to satisfy the combination criterion, we propose an online learning combinatorial strategy applicable to any number of component strategies.
In the experiment section, our proposed combinatorial strategy undergoes testing alongside other online learning multi-strategies in various designed scenarios using real data of six assets spanning three decades from CRSP. We numerically compute the proposed combinatorial strategy and demonstrate its ability to meet the investor’s objective of exceeding all component Mean-Variance strategies, as per the theoretical guarantees. It exhibits a superior growth rate and cumulative wealth compared to other online learning strategies in the experiments, albeit with a slight loss in the Sharpe ratio. Additionally, in experiments involving a large scale of component strategies, the time required for our combinatorial strategy to surpass the best one is longer than that in experiments with a small scale, consistent with theoretical analysis. Furthermore, to expedite the speed and magnitude of surpassing the cumulative wealth of the best component strategy, we propose an accelerated variant of the original combination. While this accelerated combinatorial strategy lacks a theoretical guarantee, our analysis of the experimental results demonstrates its empirical efficiency over the original strategy across any scale of component strategies.
There are certain caveats to consider when applying the proposed combinatorial strategy, particularly in tuning it for a large scale of component strategies. If there exists a single component strategy that dominates all others in terms of returns, our strategy will not exceed the cumulative wealth of the baseline one; however, the accelerated variant may achieve this. Nonetheless, such cases are often uncommon and can be addressed by adding more component strategies so that there are at least two competitive component strategies in terms of growth. In cases where the majority of the components perform poorly, the time it takes for the combinatorial strategy to exceed the best one among them will be longer, as it may experience a significant decrease before gaining enough momentum to recover. Similarly, when tuning the partition for a large scale of component strategies, it is crucial to ensure that the number of base sets is sufficient, with each base set including a majority of component strategies exhibiting similar behavior. In the context of component M-V strategies, appropriate tuning involves partitioning into evenly-sized base sets that include risk aversion levels within a small local range. In any case, acceleration can help improve the magnitude and speed of exceeding the cumulative wealth of the best component strategy.
Appendix
Proof of Proposition 2
Since $\big{(}b_{n}\big{)}\succ\big{(}\ddot{b}_{n}\big{)}$ implies $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}\land\big{(}b_{n}\big{)}\nsim\big{(}\ddot{b}_{n}\big{)}$ and by using the definition of supremum limit corresponding to the preference $\big{(}b_{n}\big{)}\succsim\big{(}\ddot{b}_{n}\big{)}$, we have:
$\forall\epsilon>0,\,\exists N_{\epsilon}\text{ such that }\sup_{n\geq N}\big{(}W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}\big{)}\big{)}<\epsilon,\,\forall N\geq N_{\epsilon}.$
Moreover, combining with the preference $\big{(}b_{n}\big{)}\nsim\big{(}\ddot{b}_{n}\big{)}$, it must result in the following:
$\forall N\geq N_{\epsilon},\,\exists n_{N}\geq N\text{ such that }W_{n_{N}}\big{(}\ddot{b}_{n_{N}}\big{)}-W_{n_{N}}\big{(}b_{n_{N}}\big{)}<0,$
which is equivalent to $S_{n_{N}}\big{(}\ddot{b}_{n_{N}}\big{)}<S_{n_{N}}\big{(}b_{n_{N}}\big{)}$, i.e., the combinatorial strategy $\left(b_{n}\right)$ is not always exceeded by the baseline strategy $\big{(}\ddot{b}_{n}\big{)}$. Otherwise, if there exists $M\geq 1$ such that $W_{n}\big{(}\ddot{b}_{n}\big{)}>W_{n}\big{(}b_{n}\big{)}$ for any $n\geq M$, then it violates the assumption $\big{(}b_{n}\big{)}\nsim\big{(}\ddot{b}_{n}\big{)}$ as:
$\epsilon>W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}\big{)}\geq 0,\,\forall n\geq\max\left\{M,N_{\epsilon}\right\},$
which implies $W_{n}\big{(}\ddot{b}_{n}\big{)}\to W_{n}\big{(}b_{n}\big{)}$.
Proof of Proposition 3
On the one hand, $\max_{\gamma}S_{n}\big{(}b_{n}^{\gamma}\big{)}\leq\max_{\lambda}S_{n}\big{(}b_{n}^{\lambda}\big{)}$ for all $n$ since $\mathcal{H}\subseteq\mathcal{\big{[}}k\big{]}$. On the other hand, since:
$\big{<}b_{n}^{\alpha^{j}},x_{n}\big{>}\geq\big{<}b_{n}^{\alpha},x_{n}\big{>},\text{ $\forall\alpha\in\mathcal{A}^{j}$,}\,j\in\left[h\right],$
then we have $S_{n}\big{(}b_{n}^{\lambda_{n}}\big{)}\leq S_{n}\big{(}b_{n}^{\gamma_{n}}\big{)}$ for all $n$ as follows:
$\max_{\lambda\in\mathcal{B}^{k}}{\displaystyle\prod_{t=1}^{n}\sum_{\alpha=1}^{k}\lambda_{\alpha}}\big{<}b_{t}^{\alpha},x_{t}\big{>}\leq\max_{\lambda\in\mathcal{B}^{k}}\prod_{t=1}^{n}\sum_{j=1}^{h}\Big{(}\sum_{\alpha\in\mathcal{A}^{j}}\lambda_{\alpha}\Big{)}\big{<}b_{t}^{\alpha^{j}},x_{t}\big{>}\leq\max_{\gamma\in\mathcal{B}^{h}}\prod_{t=1}^{n}\sum_{j=1}^{h}{\displaystyle\gamma}_{j}\big{<}b_{t}^{\alpha^{j}},x_{t}\big{>},\forall n.$
Proof of Proposition 4
By using the equality in (3.2), we have:
${\displaystyle\max_{x_{1}^{n}}\log\dfrac{S_{n}\big{(}b_{n}^{\lambda_{n}}\big{)}}{S_{n}\big{(}b_{n}\big{)}}}$ $=\max_{x_{1}^{n}}\log\dfrac{{\displaystyle\prod_{t=1}^{n}}{\displaystyle\sum_{\alpha=1}^{k}}\lambda_{n,\alpha}\big{<}b_{t}^{\alpha},x_{t}\big{>}}{{\displaystyle\int_{\mathcal{B}^{k}}}{\displaystyle\prod_{t=1}^{n}}{\displaystyle\sum_{\alpha=1}^{k}\lambda_{\alpha}}\big{<}b_{t}^{\alpha},x_{t}\big{>}\mu\left(\lambda\right)d\lambda}=\max_{x_{1}^{n}}\log\dfrac{{\displaystyle\prod_{t=1}^{n}}\big{<}\lambda_{n},r_{t}\big{>}}{{\displaystyle\int_{\mathcal{B}^{k}}}{\displaystyle\prod_{t=1}^{n}}\big{<}\lambda,r_{t}\big{>}\mu\left(\lambda\right)d\lambda},$
where the vectors $\big{(}\big{<}b_{t}^{1},x_{t}\big{>},...,\big{<}b_{t}^{k},x_{t}\big{>}\big{)}\eqqcolon\big{(}r_{t,1},...,r_{t,k}\big{)}\eqqcolon r_{t}\in\mathbb{R}_{++}^{k}$ represent the respective returns of $k$ component portfolios at time $t$. Let $r_{1}^{n}\coloneqq\big{\{}r_{1},...,r_{n}\big{\}}\in\mathbb{R}_{++}^{k\times n}$ denote the sequence of $n$ vectors of portfolios’ returns. Then, by applying the theorem of the upper bound for the worst-case of the Universal portfolio in Cover (1991); Cover and Ordentlich (1996), with replacing the sequence of assets’ returns by $r_{1}^{n}$, we obtain the upper bound with uniform density $\mu\left(\lambda\right)$ as:
$\max_{x_{1}^{n}}\log\dfrac{{\displaystyle\prod_{t=1}^{n}}\big{<}\lambda_{n},r_{t}\big{>}}{{\displaystyle\int_{\mathcal{B}^{k}}}{\displaystyle\prod_{t=1}^{n}}\big{<}\lambda,r_{t}\big{>}\mu\left(\lambda\right)d\lambda}$ $=\max_{r_{1}^{n}}\log\dfrac{{\displaystyle\prod_{t=1}^{n}}\big{<}\lambda_{n},r_{t}\big{>}}{{\displaystyle\int_{\mathcal{B}^{k}}}{\displaystyle\prod_{t=1}^{n}}\big{<}\lambda,r_{t}\big{>}\mu\left(\lambda\right)d\lambda}\leq\left(k-1\right)\log\left(n+1\right),\forall n.$
Thus, ${\displaystyle\max_{x_{1}^{n}}\left(\log S_{n}\left(b_{n}^{*}\right)-\log S_{n}\left(b_{n}\right)\right)\leq\left(k-1\right)\log\left(n+1\right)}$ for all $n$, as needed to show.
Proof of Proposition 5
On the one hand, similar to equation in (3.2), we have the following by telescoping:
$S_{n}\big{(}\hat{b}_{n}^{i}\big{)}$ $={\displaystyle{\displaystyle\prod_{t=1}^{n}\dfrac{{\displaystyle\sum_{\mathcal{A}^{i}}\big{<}{\displaystyle b_{n}^{\alpha}},x_{n}\big{>}}S_{t-1}\big{(}b_{t-1}^{\alpha}\big{)}\mu^{i}\big{(}\alpha\big{)}}{{\displaystyle\sum_{\mathcal{A}^{i}}}{\displaystyle S_{t-1}\big{(}b_{t-1}^{\alpha}\big{)}\mu^{i}\big{(}\alpha\big{)}}}}={\displaystyle\sum_{\mathcal{A}^{i}}}{\displaystyle S_{n}\left(b_{n}^{\alpha}\right)\mu^{i}\big{(}\alpha\big{)}},\,\forall i\in\left[N\right],\forall n.}$
Then, we have the following bound for every representative strategy:
${\displaystyle\log\dfrac{S_{n}\big{(}\hat{b}_{n}^{i}\big{)}}{S_{n}\big{(}b_{n}^{\alpha_{n}^{i}}\big{)}}={\displaystyle\log\dfrac{{\displaystyle\sum_{\mathcal{A}^{i}}}{\displaystyle S_{n}\left(b_{n}^{\alpha}\right)\mu^{i}\big{(}\alpha\big{)}}}{S_{n}\big{(}b_{n}^{\alpha_{n}^{i}}\big{)}}}\geq{\displaystyle\log\dfrac{S_{n}\big{(}b_{n}^{\alpha_{n}^{i}}\big{)}\mu^{i}\big{(}\alpha_{n}^{i}\big{)}}{S_{n}\big{(}b_{n}^{\alpha_{n}^{i}}\big{)}}}}$ $\geq\log\mu^{i}\big{(}\alpha_{n}^{i}\big{)},\text{$\forall i$$\in$$\left[N\right]$,$\forall n,$}$
$\geq\log\epsilon_{n},\text{$\forall i$$\in$$\left[N\right]$,$\forall n.$}$ (A.1)
On the other hand, by Proposition 4 and the property of the benchmark strategy for this context, namely $S_{n}\big{(}b_{n}^{*}\big{)}\geq\max_{i\in\left[N\right]}S_{n}\big{(}\hat{b}_{n}^{i}\big{)}$ for all $n$, we obtain the needed upper bound as follows:
$\left(N-1\right)\log\left(n+1\right)\geq\max_{x_{1}^{n}}\big{(}\log S_{n}\big{(}b_{n}^{*}\big{)}-\log S_{n}\big{(}b_{n}\big{)}\big{)}$ $\geq\max_{x_{1}^{n}}\max_{i\in\left[N\right]}\big{(}\log S_{n}\big{(}\hat{b}_{n}^{i}\big{)}-\log S_{n}\big{(}b_{n}\big{)}\big{)}$
$\geq\max_{x_{1}^{n}}\big{(}\log S_{n}\big{(}\ddot{b}_{n}\big{)}-\log S_{n}\big{(}b_{n}\big{)}\big{)}+\log\epsilon_{n},\forall n.$
The last inequality follows from (A.1) and the fact that $S_{n}\big{(}\ddot{b}_{n}\big{)}=\max_{i\in\left[N\right]}S_{n}\big{(}b_{n}^{\alpha_{n}^{i}}\big{)}$ by definition.